OpenStudy (anonymous):

How long is the minor axis for the ellipse shown below? (x+3)^2/(64) + (y-7)^2/(4) =1 A. 64 B. 16 C. 4 D. 8

OpenStudy (anonymous):

I got C?

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

did you look at the graph? minor under y ?

OpenStudy (triciaal):

(-3, 5) and (-3, 9) yes 4

OpenStudy (anonymous):

would be C)4?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

Which of the following equations represents a degenerate ellipse? 9x^2 +4y^2 = 1 x^2 y^2 ---- + ---- = 2 14 19 x^2 y^2 ----- = (-) ---- 10 18 All of the above are degenerate ellipses.

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

I got C?